OpenStudy (anonymous):

The Art Store sells three types of art pencil sets. The profit is $8.45 on watercolor pencil sets, $6.19 on pastel pencil sets and $4.62 on charcoal pencil sets. There are two branches of the Art Store. The number of sales of pencil sets for each store are shown in the table below. Set Type Sales for Westside Branch Sales for Downtown Branch watercolor 132 74 pastel 56 48 charcoal 119 126 Which store had the most profit and what was the total profit of that store? a. Westside - $1504.54 b. Westside - $2011.82 c. Downtown - $1504.54 d

OpenStudy (anonymous):

\(132\times 8.42+56\times 6.19+119\times 4.62\) is the westside profit if i am reading this correctly

OpenStudy (anonymous):

do the same think for the downtown, and compare

OpenStudy (anonymous):

*thing

OpenStudy (anonymous):

\[74\times 8.42+48\times 6.19+126\times 4.62\] for downtown

OpenStudy (anonymous):

it is a little hard to read, but i think that is right

OpenStudy (anonymous):

okay so what else do I need to do....

OpenStudy (anonymous):

compute those numbers and see which is largest

OpenStudy (anonymous):

i must have made a typo because i don't get the answer you put

OpenStudy (anonymous):

Downtown - $2011.82

OpenStudy (anonymous):

oh yeah, it is 8.45 not 8.42

OpenStudy (anonymous):

my answer was 19837.

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

Are you there.....

OpenStudy (anonymous):

back

OpenStudy (anonymous):

i get 2011.82 for the first one

OpenStudy (anonymous):

and 1504.54 for the second one

OpenStudy (anonymous):

so i would go with Westside - $2011.82